# Conformalized Late Fusion Multi-View Learning

### v1

@@ Abstract
Uncertainty quantification for multi-view learning is motivated by the increasing use of multi-view data in scientific problems. A common variant of multi-view learning is late fusion: train separate predictors on individual views and combine them after single-view predictions are available. Existing methods for uncertainty quantification for late fusion often rely on undesirable distributional assumptions for validity. Conformal prediction is one approach that avoids such distributional assumptions. However, naively applying conformal prediction to late-stage fusion pipelines often produces overly conservative and uninformative prediction regions, limiting its downstream utility. We propose a novel methodology, Multi-View Conformal Prediction (MVCP), where conformal prediction is instead performed separately on the single-view predictors and only fused subsequently. Our framework extends the standard scalar formulation of a score function to a multivariate score that produces more efficient downstream prediction regions in both classification and regression settings. We then demonstrate that such improvements can be realized in methods built atop conformalized regressors, specifically in robust predict-then-optimize pipelines.
@@ 1 Introduction
Multi-view machine learning is becoming increasingly common with the advent of multi-modal language models and conditional generative models [1, 2, 3, 4, 5]. In addition, such techniques are being increasingly used in more safety-critical settings, such as medicine [6, 7, 8, 9] and robotics [10, 11, 12, 13]. These use cases motivate the development of uncertainty quantification methods for multi-view learning.
The most common division of multi-view learning is along the axis of when fusion of the multiple views is performed. We can either learn a fused data representation or we can fuse the predictions from the separate views. The latter approach is referred to as “late fusion”. It is desirable in limited resource settings, where the separately trained models can be combined in a light-weight fashion for downstream tasks [14, 15, 16, 17]. Increasing interest is now being placed on quantifying uncertainty for such models [18, 19, 20, 21, 22]. Such methods, however, rely on distributional assumptions to provide coverage guarantees or are computationally intractable, limiting their practical utility. These limitations, therefore, demand a more universal, yet informative, uncertainty quantification method.
@@
One such method for performing distribution-free uncertainty quantification is conformal prediction, which provides a principled framework for producing distribution-free prediction regions with marginal frequentist coverage guarantees [23, 24]. By using conformal prediction on a user-defined score function $s(x,y)$ and obtaining an empirical $1-\alpha$ quantile $\widehat{q}(\alpha)$ of $s(x,y)$ over a calibration set $\mathcal{D}_{C}$, prediction regions $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}(\alpha)\}$ attain marginal coverage guarantees. While calibration is guaranteed from this procedure, predictive efficiency, quantified as the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
We extend the conformal prediction line of work to enable its practical use in late-stage multi-view prediction problems. In particular, we propose a multivariate extension to the score function notion introduced in traditional conformal prediction, with a corresponding extension of quantiles to “quantile envelopes.” We demonstrate that this formulation retains the desired distribution-free coverage guarantees typical of standard conformal prediction and that the resulting prediction regions can be used efficiently in both classification and regression settings. Our contributions are:
@@
Figure 1: MVCP provides a principled extension to the standard conformal prediction pipeline by leveraging ideas from higher-dimensional quantile regression to define quantile envelopes $\widehat{\mathcal{Q}}$ instead of purely scalar quantiles $\widehat{q}$. It does so by evaluating a collection of score functions (here $s_{1}$ and $s_{2}$) over the calibration dataset to define $\mathcal{S}$, finding quantiles $\{\widehat{q}_{m}\}$ over a set of projection directions $\{u_{m}\}$, and taking $\widehat{\mathcal{Q}}$ to be the intersection of the resulting half-planes $H(u_{m},\widehat{q}_{m})$. These quantile envelopes result in more informative prediction regions that can be used in downstream tasks.
  • Providing a multivariate extension to conformal prediction score functions (MVCP) that retains coverage guarantees and can be applied to late-stage multi-view prediction problems.
@@
  • Demonstrating how the resulting prediction regions can be efficiently leveraged for downstream classification and regression tasks.
  • Demonstrating the empirical performance of the MVCP framework across a suite of tasks.
@@ 2.1 Conformal Prediction
Coverage guarantees of uncertainty quantification methods generally rely on distributional assumptions, often via asymptotics or explicit model specification. To alleviate the need for such restrictive assumptions, interest in finite-sample, distribution-free uncertainty quantification methods has risen. Conformal prediction is one such method [23, 24]. Due to its lack of structural assumptions, conformal prediction pairs well with the increasingly common black-box style of modeling.
@@
Conformal prediction serves as a wrapper around such predictors, producing prediction regions $\mathcal{C}(x)$ that have formal guarantees of the form $\mathcal{P}_{X,Y}(Y\notin\mathcal{C}(X))\leq\alpha$ for some prespecified level $\alpha$. To achieve this, “split conformal” partitions the dataset $\mathcal{D}$ into a training set $\mathcal{D}_{T}$ and a calibration set $\mathcal{D}_{C}$. The former plays the role traditionally played by $\mathcal{D}$, namely serving as data used to fit $\widehat{f}$. The latter is used much in the way a validation set is typically employed in practice. Formally, users of conformal prediction must design a “score function” $s(x,y)$, which should quantify “test error,” often in a domain-specific manner. For instance, a simple score function for a regression setting would be $s(x,y)=\|\hat{f}(x)-y\|$. This score function is then evaluated across the calibration set to define $\mathcal{S}=\{s(x,y)\mid(x,y)\in\mathcal{D}_{C}\}$. For a desired coverage of $1-\alpha$, we then take $\widehat{q}$ to be the $\lceil(|\mathcal{D}_{C}|+1)(1-\alpha)\rceil/|\mathcal{D}_{C}|$ quantile of $\mathcal{S}$, with which prediction regions for future test queries $x$ can be defined as $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}\}$. Under the exchangeability of the score of a test point $s(x^{\prime},y^{\prime})$ with $\mathcal{S}$, we have the desired finite-sample probabilistic guarantee that $1-\alpha\leq\mathcal{P}_{X,Y}(Y\in\mathcal{C}(X))$.
@@
While this guarantee holds for any $s(x,y)$, the informativeness of the resulting prediction regions, formally defined as the inverse expected Lebesgue measure across $X$, i.e. $\left(\mathbb{E}[\mathcal{L}(\mathcal{C}(X))]\right)^{-1}$, is intimately tied to its specification [24]. Thus, much of the practical challenge of conformal prediction relates to choosing a score function that retains coverage while minimizing prediction region size.
@@ 2.2 Predict-Then-Optimize
Prediction regions produced from conformalized predictors are often leveraged differently across classification and regression settings. In the case of classification, “regions” simply constitute a subset of the label space, making its direct use by end users straightforward, as demonstrated in works such as [25]. In high-dimensional regression settings, however, prediction regions become harder to use directly; for this reason, recent works have started shifting focus to using them in their implicit forms.
@@
While the predicted $\widehat{c}$ is often simply trusted in the downstream optimization, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to highly suboptimal decision-making. For this reason, recent interest has been placed on studying a “robust” alternative to this formulation [27, 28, 29]. Following this line of work, [26] proposed studying the following formulation: $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with the uncertainty region $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$. The primary consequence of doing so was having a probabilistic guarantee on the suboptimality gap. That is, denoting the suboptimality gap by $\Delta(x,c):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})-\min_{w}f(w,c)$, where $c$ is the true parameter corresponding to $x$, they established that leveraging conformal uncertainty regions guarantees $\mathcal{P}_{X,C}\left(0\leq\Delta(X,C)\leq L\mathrm{\ diam}(\mathcal{C}(X))\right)\geq 1-\alpha$ if $f(w,c)$ is an objective function that is convex-concave and $L$-Lipschitz in $c$ for any fixed $w$.
2.3 Multi-View Learning
We provide a brief overview of multi-view learning herein, deferring to recent surveys of the field for a more comprehensive review thereof [30]. In general, the focus of multi-view machine learning is to solve prediction problems when different, often non-overlapping, representations of the same underlying “object” are observed. Methods in this space bifurcate between those defined together with some downstream supervised task and those seeking to find an abstract “fused” representation that lends itself naturally for arbitrary downstream use cases. We focus herein on the former category. For instance, one may wish to classify a captioned image, where the image and caption represent two distinct “views” of the same underlying abstract object in this case.
Formally, therefore, we consider the multi-view prediction problem as seeking a map $\mathcal{X}\rightarrow\mathcal{Y}$ with $\mathcal{X}=\mathcal{X}_{1}\cup...\cup\mathcal{X}_{K}$. As previously mentioned, fusion approaches are generally either early-stage or late-stage. In the former, we seek to learn some common, often abstract, space into which these separate views can be mapped for downstream prediction, namely seeking a map $g_{\phi}$ that maps $g_{\phi}:\mathcal{X}_{1}\times...\times\mathcal{X}_{K}\rightarrow\mathcal{Z}$. Such a fusion map is then assumed to be usable for the downstream prediction task, namely through composition with $f_{\theta}:\mathcal{Z}\rightarrow\mathcal{Y}$ such that the overall prediction model can be defined as $f\circ g$. Traditionally, $g$ was hand-crafted, leveraging domain-specific insights. However, increasingly, $f$ and $g$ are simply subsets of a neural network trained end-to-end via gradient descent, such as in [31, 32, 33, 34, 35, 36].
As discussed previously, this fusion approach is not universally desirable in all circumstances though, such as when models for individual views have already been curated after computationally intensive training and hyperparameter tuning [14, 15, 16, 17]. In such cases, late-stage fusion approaches can be preferred, where separate predictors $f_{i}:\mathcal{X}_{i}\rightarrow\mathcal{Y}$ have been trained and are fused via a post-hoc algorithm $\mathcal{F}:\mathcal{Y}\times...\times\mathcal{Y}\rightarrow\mathcal{Y}$. Approaches to this include decision averaging [37, 38], decision weighting [18, 39, 40], and decision voting [41]. We focus herein on such late-stage fusion.
@@ 2.4 Related Works
Work in the conformal prediction literature has yet to leverage the generalized notions of quantiles considered herein. Generalizations of quantiles, while introduced first in the context of conformal prediction herein, have a long history in traditional statistics, with recent surveys on approaches available in [42, 43]. We briefly summarize the relevant pieces of this literature below.
Unlike univariate data, multivariate data does not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering of points in higher dimensional spaces. To account for this, one class of multivariate quantiles, known as “direction quantiles,” prescribes directions along which data points can be compared [44, 45, 46]. That is, for some direction $u\in\mathcal{S}^{n-1}$, points $x\in\mathbb{R}^{n}$ are projected, after which the standard notion of a quantile can be leveraged.
The notion of the $\alpha$ quantile for a random variable $X\in\mathcal{X}$ clearly then depends on this choice of direction $u$, which we denote by $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$. When there is no ambiguity, we just denote it as $Q(\alpha,u)$. In cases where domain knowledge assigns an unambiguous notion of multivariate ordering, such knowledge can be encoded into the choice of $u$. In other cases, however, a more agnostic approach is typically employed: rather than choosing any particular direction, all directions are considered simultaneously. Formally, for any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$. To extend this to consider all directions simultaneously, we then define the quantile envelope to be
$D(\alpha)=\bigcap_{u\in\mathcal{S}^{n-1}}H(u,Q(\alpha,u)).$ (1)
To the best of our knowledge, the only work leveraging conformal prediction in the context of multivariate quantiles is [47]. This work, however, focuses on a fundamentally different problem than that considered herein. In particular, they focused on leveraging conformal prediction as a correction step to the output of directional quantile regression with the intended use case being in constructing calibrated prediction regions for multivariate regression problems. This work, therefore, was a natural extension to the now classical application of conformal prediction for the correction of univariate quantile regression. Our proposed method, however, does the converse: it leverages directional quantiles within the score space to extend conformal prediction itself.

### v2

# Conformal Prediction for Ensembles: Improving Efficiency via Score-Based Aggregation

@@ Abstract
Distribution-free uncertainty estimation for ensemble methods is increasingly desirable due to the widening deployment of multi-modal black-box predictive models. Conformal prediction is one approach that avoids such distributional assumptions. Methods for conformal aggregation have in turn been proposed for ensembled prediction, where the prediction regions of individual models are merged as to retain coverage guarantees while minimizing conservatism. Merging the prediction regions directly, however, sacrifices structures present in the conformal scores that can further reduce conservatism. We, therefore, propose a novel framework that extends the standard scalar formulation of a score function to a multivariate score that produces more efficient prediction regions. We then demonstrate that such a framework can be efficiently leveraged in both classification and predict-then-optimize regression settings downstream and empirically show the advantage over alternate conformal aggregation methods.
@@ 1 Introduction
Ensemble methods are an oft-used class of statistical modeling techniques due to their ability to reduce variance or improve predictive accuracy [1, 2, 3]. Such methods are increasingly being coupled with complex, black-box models, such as in multi-modal language models [4, 5, 6, 7, 8]. Couplings of this sort are seeing ever-widening deployment in safety-critical settings, such as medicine [9, 10, 11] and robotics [12, 13, 14].
Increasing interest is, therefore, now being placed on quantifying uncertainty for such models [15, 16, 17, 18, 19]. Towards this end, methods of uncertainty quantification have arisen, such as deep ensembles and committee estimation [20, 21, 22]. Such methods, however, sacrifice generality with the imposition of distributional assumptions, motivating the need for distribution-free uncertainty quantification for ensemble methods.
@@
One method for performing distribution-free uncertainty quantification is conformal prediction, which provides a principled framework for producing distribution-free prediction regions with marginal frequentist coverage guarantees [23, 24]. By using conformal prediction on a user-defined score function, prediction regions attain marginal coverage guarantees. While calibration is guaranteed from this procedure, predictive efficiency, quantified as the size of the resulting prediction regions, can be unboundedly large for poorly chosen score functions.
As a result, methods have arisen to perform conformal model aggregation, which both provide uncertainty estimates of the ensembled predictions and do so in ways as to minimize the prediction region size [25, 26, 27, 28, 29]. While such approaches succeed in reducing the prediction region size over naive aggregation, they all aggregate the separately conformalized prediction regions of the predictors in the ensemble. In doing so, they forgo the possibility of automatically leveraging shared structure amongst the scores of the individual predictors, resulting in overly conservative prediction regions.
@@
Figure 1: CSA provides a principled extension to the standard conformal prediction pipeline by leveraging ideas from higher-dimensional quantile regression to define quantile envelopes $\widehat{\mathcal{Q}}$ instead of scalar quantiles $\widehat{q}$. It does so by evaluating a collection of score functions (here $s_{1}$ and $s_{2}$) over the calibration dataset to define $\mathcal{S}$, finding quantiles $\{\widehat{q}_{m}\}$ over a set of projection directions $\{u_{m}\}$, and taking $\widehat{\mathcal{Q}}$ to be the intersection of the resulting half-planes $H(u_{m},\widehat{q}_{m})$. These quantile envelopes result in more informative prediction regions that can be used in downstream tasks.
We instead propose to perform aggregation in score space by extending traditional conformal prediction to consider a multivariate score function and defining prediction regions using “quantile envelopes” in place of scalar quantiles. Doing so enables efficient, data-driven, automated conformal model aggregation. We demonstrate that this formulation retains the desired distribution-free coverage guarantees typical of standard conformal prediction and that the resulting prediction regions can be used efficiently in both classification and regression settings. Our contributions, thus, are:
  • Providing a multivariate extension to conformal prediction, dubbed “conformal score aggregation” (CSA), that leverages quantile envelopes to enable data-driven, informative uncertainty estimation for model ensembles while retaining coverage guarantees.
@@
  • Demonstrating how the prediction regions resulting from CSA can be efficiently leveraged in downstream predict-then-optimize regression tasks.
  • Demonstrating the empirical improvement of the CSA framework over alternate conformal aggregation strategies across classification and regression settings.
@@ 2.1 Conformal Prediction
Coverage guarantees of uncertainty quantification methods generally rely on distributional assumptions, often via asymptotics or explicit specification. To alleviate the need for such restrictive assumptions, interest in finite-sample, distribution-free uncertainty quantification methods has risen. Conformal prediction is one such method [23, 24]. Due to its lack of structural assumptions, conformal prediction pairs well with the increasingly common black-box style of modeling.
@@
Conformal prediction serves as a wrapper around such predictors, producing prediction regions $\mathcal{C}(x)$ that have formal guarantees of the form $\mathcal{P}_{X,Y}(Y\notin\mathcal{C}(X))\leq\alpha$ for some prespecified level $\alpha$. To achieve this, “split conformal” partitions the dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ into a training set $\mathcal{D}_{T}$ and a calibration set $\mathcal{D}_{C}$. The former serves as the data used to fit $\widehat{f}$. Users of conformal prediction must then design a “score function” $s(x,y)$, which should quantify “test error”, often in a domain-specific manner. For instance, a simple score function for a regression setting would be $s(x,y)=\|\hat{f}(x)-y\|$. This score function is then evaluated across the calibration set to define $\mathcal{S}_{C}=\{s(x,y)\mid(x,y)\in\mathcal{D}_{C}\}$. For a desired coverage of $1-\alpha$, we then take $\widehat{q}$ to be the $\lceil(|\mathcal{D}_{C}|+1)(1-\alpha)\rceil/|\mathcal{D}_{C}|$ quantile of $\mathcal{S}_{C}$, with which prediction regions for future test queries $x$ can be defined as $\mathcal{C}(x)=\{y\mid s(x,y)\leq\widehat{q}\}$. Under the exchangeability of the score of a test point $s(X^{\prime},Y^{\prime})$ with $\mathcal{S}_{C}$, we have the desired finite-sample probabilistic guarantee that $1-\alpha\leq\mathcal{P}_{X^{\prime},Y^{\prime}}(Y^{\prime}\in\mathcal{C}(X^{\prime}))$.
@@
While this guarantee holds for any $s(x,y)$, the informativeness of the resulting prediction regions, quantified as the inverse expected Lebesgue measure across $X$, i.e. $\left(\mathbb{E}[\mathcal{L}(\mathcal{C}(X))]\right)^{-1}$, is intimately tied to its specification [24]. Thus, much of the challenge of conformal prediction relates to choosing a score function that retains coverage while minimizing region size.
2.2 Quantile Envelopes
Generalizations of quantiles have a long history in statistics [30, 31]. Unlike univariate data, multivariate data does not lend itself to an unambiguous definition of a quantile, as there is no canonical ordering in higher dimensional spaces. To account for this, one class of multivariate quantiles, known as “direction quantiles,” prescribes directions along which data points can be compared [32, 33, 34]. That is, points $x\in\mathbb{R}^{n}$ are projected along some direction $u\in\mathcal{S}^{n-1}$, after which the standard notion of a quantile can be leveraged.
The notion of the $\alpha$ quantile for a random variable $X\in\mathcal{X}$ clearly then depends on this choice of direction $u$, which we denote by $Q(X,\alpha,u)=\inf\{q\in\mathbb{R}:\mathcal{P}(u^{\top}X\leq q)\geq\alpha\}$. When there is no ambiguity, we just denote it as $Q(\alpha,u)$. In cases where domain knowledge assigns an unambiguous notion of multivariate ordering, such knowledge can be encoded into the choice of $u$. In other cases, however, a more agnostic approach is employed: rather than choosing any particular direction, all directions are considered simultaneously. For any given $u$, notice the choice of quantile defines a corresponding halfplane $H(u,Q(\alpha,u))=\left\{x\in\mathcal{X}:u^{\top}x\leq Q(\alpha,u)\right\}$. To consider all directions simultaneously, the quantile envelope is then defined to be
$D(\alpha)=\bigcap_{u\in\mathcal{S}^{n-1}}H(u,Q(\alpha,u)).$ (1)
Notably, while each individual $H(u,Q(\alpha,u))$ captures $1-\alpha$ of the points, $D(\alpha)$ does not, as it is the intersection thereof and hence captures $<1-\alpha$ of the mass. If $1-\alpha$ combined coverage is sought, a correction, such as Bonferroni adjustment, is used for the individual planes.
@@ 2.3 Predict-Then-Optimize
Prediction regions produced from conformalized predictors are often leveraged differently across classification and regression settings. In the case of classification, “regions” simply constitute a subset of the label space, making their direct use by end users straightforward, as demonstrated in works such as [35]. In high-dimensional regression settings, however, prediction regions become harder to use directly; for this reason, recent works have started shifting focus to using them in their implicit forms.
@@
While the predicted $\widehat{c}$ is often simply trusted in the downstream optimization, this approach is inappropriate in risk-sensitive settings, where misspecification of the map $g:\mathcal{X}\rightarrow\mathcal{C}$ could lead to suboptimal decision-making. For this reason, recent interest has been placed on studying a “robust” alternative to this formulation [37, 38, 39]. Following this line of work, [36] proposed studying $w^{*}(x):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})$, with $\mathcal{C}(x)$ being produced by conformalizing the predictor $g$. The primary consequence of doing so was having a probabilistic guarantee on the suboptimality gap. That is, denoting the suboptimality gap by $\Delta(x,c):=\min_{w}\max_{\widehat{c}\in\mathcal{C}(x)}f(w,\widehat{c})-\min_{w}f(w,c)$, where $c$ is the true parameter corresponding to $x$, they established that leveraging conformal uncertainty regions guarantees $\mathcal{P}_{X,C}\left(0\leq\Delta(X,C)\leq L\mathrm{\ diam}(\mathcal{C}(X))\right)\geq 1-\alpha$ if $f(w,c)$ is an objective function that is convex-concave and $L$-Lipschitz in $c$ for any fixed $w$.
@@ 2.4 Related Works
Ensemble methods consist of $K$ predictors $f_{k}:\mathcal{X}_{k}\rightarrow\mathcal{Y}$; notably, such predictors need not map from the same set of covariates. A naive approach for uncertainty quantification would then be to conformalize the ensembled predictor. That is, for an ensembling algorithm $\mathcal{F}:\mathcal{Y}^{K}\rightarrow\mathcal{Y}$, a score function $s(\mathcal{F}(f_{1}(x),...,f_{K}(x)),y)$ would be defined. Denoting the $\lceil(N_{\mathcal{C}}+1)(1-\alpha)\rceil/N_{\mathcal{C}}$ quantile of the score distribution over $\mathcal{D}_{C}$ as $\widehat{q}(\alpha)$, $\mathcal{C}(x)=\{y:s(x,y)\leq\widehat{q}(\alpha)\}$ would then be marginally calibrated.
Such an approach, however, lacks some desirable properties. In particular, prediction regions $\mathcal{C}(x)$ should have the quality that, if a particular predictor has less uncertainty in its predictions, as is frequently true of ensemble settings where the predictors span multiple input data modalities, upon routing to that predictor, the corresponding size of the prediction region should be smaller than if it had been routed to a different predictor. While the naive approach does, in principle, support this property, it ultimately relies on defining an uncertainty-aware ensembling algorithm $\mathcal{F}$. In its typical form, however, $\mathcal{F}$ simply takes point predictions $f_{1}(x),...,f_{K}(x)$ in as input, meaning any uncertainty-awareness would need to be baked in a priori into the definition of $\mathcal{F}$ through domain knowledge of the uncertainties of the predictors $f_{1},...,f_{K}$, which can seldom be specified precisely, sacrificing the predictive efficiency of $\mathcal{C}(x)$.
Conformal model aggregation, thus, seeks to mitigate these deficiencies by aggregating the prediction regions $\mathcal{C}_{1}(x),...,\mathcal{C}_{K}(x)$ rather than the individual point predictions [25, 27, 28, 29]. While there are several methods in this vein, they can be categorized into one of two general approaches. The first line of work seeks to perform model selection, in which a single conformal predictor is selected $\mathcal{C}_{k^{*}}$, typically based on the criterion of minimizing region size $k^{*}:=\operatorname*{arg\,min}_{k}\mathbb{E}[\mathcal{L}(\mathcal{C}_{k}(X))]$ [27, 28]. Generally, however, methods leveraging the full collection of predictors produce less conservative regions [25, 29]. Such works aggregate the individual prediction regions into a final region by defining
$\mathcal{C}(x):=\left\{y\mid\sum_{k=1}^{K}w_{k}\mathbbm{1}[y\in\mathcal{C}_{k}(x)]\geq\widehat{a}\right\}$ (2)
for weights $\{w_{k}\}\in[0,1]$ such that $\sum_{k=1}^{K}w_{k}=1$ and a threshold $\widehat{a}$. Methods then differ in the procedure by which $\{w_{k}\}$ and $\widehat{a}$ are prescribed, several of which were prescribed in [29], whose detailed presentation is deferred to Appendix F for space reasons. We note that the methods of [25] are designed for a different setting than that considered herein, namely that in which conformal coverage is sought adaptively over online data streams.
In this vein, [40] is a recent work that has similarly proposed a vector-score extension as that discussed herein, in which candidate weight vectors $\{w_{m}\}\in\mathbb{R}^{K}$ are searched over for score aggregation. That is, a vector $s(x):=(s_{1}(x,y),...,s_{K}(x,y))\in\mathbb{R}^{K}$ of scores $s_{k}(x,y)$ corresponding to each predictor $f_{k}(x)$ is predicted and its aggregate prediction region defined on the projection $\langle w_{m^{*}},s\rangle$ for $w_{m^{*}}$ the weight resulting in the smallest prediction region. This method, however, has two shortcomings addressed herein. The first is that their method can only be applied in classification settings, whereas our method can be leveraged across both regression and classification problems. The second is that their approach only uses a single weighted projection in the end, resulting in suboptimal aggregation and, therefore, conservative prediction regions.